Under the same conditions of temperature and pressure, does hydrogen iodide (HI) or

ammonia (NH3) effuse faster? Calculate the relative rates at which they effuse.



What is the effusion rate of both HI and NH3?
1
Expert's answer
2020-09-17T08:21:10-0400

According to Graham's law of effusion, the rate of effusion of a gas is inversely proportional to the square root of its molecular weight. Therefore, the gas with smaller molecular weight (NH3) effuses faster than the one with greater molecular weight (HI). For two gases:

"r_1/r_2=\sqrt{M_2 /M_1}"

"r_{NH_3}/r_{HI} = \sqrt{M_{HI}/M_{NH_3}} = \sqrt{127.91/17.3}=7.39"


The effusion rate of NH3 is 7.39 times higher than the one of HI.